Task 4 make your own experiment and trial .choose your own random variable X and compute for its mean , variance and standard deviation





1
Expert's answer
2022-03-17T07:19:25-0400

Solution

Let "X" represent the marks of "7" students in a test


"X=66,73,59,62,69,65,70"


(a) mean "\bar X=\dfrac{\sum X_i}n"


"\bar X=\frac{66+73+59+62+69+65+70}{7}"


"\bar X=\dfrac{464}7=66.288"


(b) Variance


"~~~~~~~\def\arraystretch{1.5}\begin{array}{c:c:c}X_i
&X_i -\bar X& (X_i -\bar X)^2\\\hline59&-7.288&53.115\\\hdashline62&-4.288&18.387\\\hdashline65&-1.288&1.659\\\hdashline66&-0.288&0.083\\\hdashline69&2.712&7.355\\\hdashline70&3.712&13.779\\\hdashline73&6.712&45.051\\\hline\sum&&139.429\\\hline\end{array}"


"\sigma^2=\dfrac{139.429}7=19.92"


(c) Standard deviation


"\sigma=\sqrt{variance}"


"\sigma=\sqrt{19.92}"


"\sigma=4.463"